# TRIPP: A General Purpose Data Pipeline for Astronomical Image Processing

### v1

@@ Abstract
We present the TRansient Image Processing Pipeline (TRIPP), a transient and variable source detection pipeline that employs both difference imaging and light curve analysis techniques for astronomical data. Additionally, we demonstrate TRIPP’s rapid analysis capability by detecting transient candidates in near-real time. TRIPP was tested using image data of the supernova SN2023ixf and from the Local Galactic Transient Survey (LGTS; LGTS) collected by the Las Cumbres Observatory’s (LCO) network of 0.4 m telescopes. To verify the methods employed by TRIPP, we compare our results to published findings on the photometry of SN2023ixf. Additionally, we report the ability of TRIPP to detect transient signals from optical Search for Extra Terrestrial Intelligence (SETI) sources.
@@ 2.1 LGTS Data and Optical SETI
The majority of real data used to test TRIPP is from our Local Galactic Transient Survey (LGTS) using Las Cumbres Observatory Global Telescope network (LCOGT) 0.4 m telescopes to gather 6 megapixel images of the Andromeda Galaxy and the Magellanic Clouds. In LGTS, we explore the results of Lubin (2016) to demonstrate how our general-purpose transient pipeline can utilize LGTS images for the Search for ExtraTerrestrial Intelligence (SETI). Based on our understanding of terrestrial technologies and future capabilities, SETI has historically focused on radio frequencies, which were well understood as means of communication when efforts began in the 1960s (see Tarter (2001)). However, recent advances in laser technology suggest that lasers may be an effective means of interstellar communication. Since few SETI searches have examined the optical wavelengths available with LCOGT telescopes, this is a natural extension of SETI efforts (Price et al. (2020), Hippke (2018), Lubin (2016)). Our findings in LGTS demonstrated that TRIPP could feasibly detect optical SETI signals in LGTS data with LCOGT 0.4m telescopes. Therefore, by processing LGTS data, we aim to validate TRIPP’s performance at relatively low signal-to-noise (due to the surface brightness of target galaxies) while making a concerted effort towards SETI.
@@ 6 Conclusion
We are presently utilizing TRIPP to detect Milky Way transients obscured by the high surface brightness of the Andromeda galaxy along with bright transients from the Andromeda galaxy. These results validate pipeline performance at a low signal-to-noise. In addition, as TRIPP is capable of detecting optical SETI transients from civilizations with lasers that we can construct within the century (see LGTS, Lubin (2016)), Andromeda is an ideal target due to the large number of possible messenger civilizations and relative proximity.

### v2

@@ Abstract
We present the TRansient Image Processing Pipeline (TRIPP), a transient and variable source detection pipeline that employs both difference imaging and light curve analysis techniques for astronomical data. Additionally, we demonstrate TRIPP’s rapid analysis capability by detecting transient candidates in near-real time. TRIPP was tested using image data of the supernova SN2023ixf and from the Local Galactic Transient Survey (LGTS; Thomas et al. (2025)) collected by the Las Cumbres Observatory’s (LCO) network of 0.4 m telescopes. To verify the methods employed by TRIPP, we compare our results to published findings on the photometry of SN2023ixf. Additionally, we report the ability of TRIPP to detect transient signals from optical Search for Extra Terrestrial Intelligence (SETI) sources.
@@ 2.1 LGTS Data and Optical SETI
The majority of real data used to test TRIPP is from our Local Galactic Transient Survey (LGTS) using Las Cumbres Observatory Global Telescope network (LCOGT) 0.4 m telescopes to gather 6 megapixel images of the Andromeda Galaxy and the Magellanic Clouds. In Thomas et al. (2025), we explore the results of Lubin (2016) to demonstrate how our general-purpose transient pipeline can utilize LGTS images for the Search for ExtraTerrestrial Intelligence (SETI). Based on our understanding of terrestrial technologies and future capabilities, SETI has historically focused on radio frequencies, which were well understood as means of communication when efforts began in the 1960s (see Tarter (2001)). However, recent advances in laser technology suggest that lasers may be an effective means of interstellar communication. Since few SETI searches have examined the optical wavelengths available with LCOGT telescopes, this is a natural extension of SETI efforts (Price et al. (2020), Hippke (2018), Lubin (2016)). Our findings in Thomas et al. (2025) demonstrated that TRIPP could feasibly detect optical SETI signals in LGTS data with LCOGT 0.4m telescopes. Therefore, by processing LGTS data, we aim to validate TRIPP’s performance at relatively low signal-to-noise (due to the surface brightness of target galaxies) while making a concerted effort towards SETI.
@@ 6 Conclusion
We are presently utilizing TRIPP to detect Milky Way transients obscured by the high surface brightness of the Andromeda galaxy along with bright transients from the Andromeda galaxy. These results validate pipeline performance at a low signal-to-noise. In addition, as TRIPP is capable of detecting optical SETI transients from civilizations with lasers that we can construct within the century (see Thomas et al. (2025), Lubin (2016)), Andromeda is an ideal target due to the large number of possible messenger civilizations and relative proximity.